• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name W. Boeglin

    Log entry time 16:08:59 on November 8,2002

    Entry number 89860

    keyword=Shift Summary

    Continue taking data for Q3D_i40

    Run Start End Type
    3077 8:35 Production, 1M (many trips, every 4min)
    3078 8:36 9:15 177K, no beam for 1/2 hour
    3079 9:17 9:36 195K
    3080 9:45 no beam for the last 10 min.
    3081 12:54 14:07 1M Production

    8:50 MCC called, vacuum convectron problem in arc1, no time estimate so far
    9:15 " problem seems to be fixed -> expect beam back
    9:35 Cavity problems -> no beam, no time estimate yet
    9:45 beam seems to come back, is very unstable
    10:05 MCC announced beam interruptions to work on increasing the beam current for Hall C as agreed upon in this mornings meeting. This is expected to take a maximum of 1 hour.
    11:10 talked to PD about getteing the beam back. I agreed to wait until noon so that they can optimize the orbits for Hall C

    12:15 MCC announced that they will start to return beam
    12:50 beam starts to return (and trips immediately)
    12:54 usable beam back
    14:10 beam turned off for re-tuning due to too much loss in the machine.

    15:35 MCC called and announced that they need at least another 1/2 hour.
    16:00 we are still waiting for beam